### 11 Jan 2024 v1

@@ 1 Introduction
Additionally, accurately modeling both the inter-dependencies between context and interlocutors, as well as the intra-dependencies within speakers in dialogues, remains a significant challenge. Previous homogeneous models, such as the graph attention network Veličković et al. (2017) Brody et al. (2021), did not consider the variations in link types. Heterogeneous models like relational graph networks (RGCN) employ distinct relation types to model various dependencies. Yet, they utilize shared coefficients across all relation types, which may fail to capture the unique attributes of each relation type, as shown in Fig 1. To address this issue, we propose a method to independently model heterogeneous conversational interactions, capturing both contextual influences and inherent personality traits. Our main contributions are as follows:
@@
  • We propose a heterogeneous conversational graph network (HC-GNN) to independently model both the interdependencies among interlocutors, as well as the intra-dependencies within the speaker in dialogues.
@@ 4.2 Heterogeneous Conversational Graph Neural Network
We map each utterance in the dialogue into embeddings, as presented in Section 4.2. Subsequently, we describe heterogeneous conversation modeling in Section 4.2, followed by an explanation of heterogeneous conversational graph network feature encoding and fusion in Section 4.2.
@@
To explicitly model the interaction between speakers, we independently model the intra-dependency (innate personality) and inter-dependency (acquired personality which is influenced by the interlocutor), as shown in Fig m 2. We introduce the following notation: we denote directed and labeled multi-graphs as $\mathcal{G}_{r}=(\mathcal{V}_{r},\mathcal{E}_{r},r)$ with nodes $\vartheta_{r,i}\in\mathcal{V}_{r}$ and labeled edges (relations) $(\vartheta_{r,i},r,\vartheta_{r,j})\in\mathcal{E}_{r}$, where $r\in\mathcal{R}$ represents one of the conversation relation types {$spk$ $a$ normal-$\rightarrow$ $spk$ $a$, $spk$ $b$ normal-$\rightarrow$ $spk$ $b$, $spk$ $b$ normal-$\rightarrow$ $spk$ $a$, $spk$ $a$ normal-$\rightarrow$ $spk$ $b$} .
@@ 6.1 Monologue
2. Speaker Choice. We compare generating synthetic dialogue from the same speaker or two different speakers. We observe using dialogues from different speakers not only enables continuous data distribution as shown in Fig 3 but also showed much higher averaged accuracy ($61.2\%$ vs $57.3\%$) and balanced accuracy ($60.4\%$ vs $58.1\%$), which demonstrates that speaker variety is more crucial than the number of conversations for personality recognition within this dataset.

### 08 Mar 2024 v2

@@ 1 Introduction
Additionally, accurately modeling both the inter-dependencies between context and interlocutors, as well as the intra-dependencies within speakers in dialogues, remains a significant challenge. Previous homogeneous models, such as the graph attention network Veličković et al. (2017) Brody et al. (2021), did not consider the variations in link types. Heterogeneous models like relational graph convolution networks (RGCN) employ distinct relation types to model various dependencies. Yet, they utilize shared coefficients across all relation types, which may fail to capture the unique attributes of each relation type, as shown in Fig 1. To address this issue, we propose a method to independently model heterogeneous conversational interactions, capturing both contextual influences and inherent personality traits. Our main contributions are as follows:
@@
  • We propose a heterogeneous conversational graph neural network (HC-GNN) to independently model both the interdependencies among interlocutors, as well as the intra-dependencies within the speaker in dialogues.
@@ 4.2 Heterogeneous Conversational Graph Neural Network
We map each utterance in the dialogue into embeddings, as presented in Section 4.2. Subsequently, we describe heterogeneous conversation modeling in Section 4.2, followed by an explanation of heterogeneous conversational graph neural network feature encoding and fusion in Section 4.2.
@@
To explicitly model the interaction between speakers, we independently model the intra-dependency (innate personality) and inter-dependency (acquired personality which is influenced by the interlocutor), as shown in Fig. 2. We introduce the following notation: we denote directed and labeled multi-graphs as $\mathcal{G}_{r}=(\mathcal{V}_{r},\mathcal{E}_{r},r)$ with nodes $\vartheta_{r,i}\in\mathcal{V}_{r}$ and labeled edges (relations) $(\vartheta_{r,i},r,\vartheta_{r,j})\in\mathcal{E}_{r}$, where $r\in\mathcal{R}$ represents one of the conversation relation types {$spk$ $a$ normal-$\rightarrow$ $spk$ $a$, $spk$ $b$ normal-$\rightarrow$ $spk$ $b$, $spk$ $b$ normal-$\rightarrow$ $spk$ $a$, $spk$ $a$ normal-$\rightarrow$ $spk$ $b$} .
@@ 6.1 Monologue
2. Speaker Choice. We compare generating synthetic dialogue from the same speaker or two different speakers. We observe using dialogues from different speakers not only enables continuous data distribution as shown in Fig. 3 but also showed much higher averaged accuracy ($61.2\%$ vs $57.3\%$) and balanced accuracy ($60.4\%$ vs $58.1\%$), which demonstrates that speaker variety is more crucial than the number of conversations for personality recognition within this dataset.